1.35g of aluminum powder and 8.0 g of copper (II) oxide were used in excess and heated vigorously in the laboratory to produce aluminum oxide and copper.


Calculate the mass of copper (ii) oxide not used in the reaction.


1
Expert's answer
2021-10-08T02:09:15-0400

2Al + 3CuO → 3Cu + Al2O3

M(Al) = 26.98 g/mol

n(Al) "= \frac{1.35}{26.98} = 0.05 \;mol"

According to the reaction:

n(CuO) "= \frac{3}{2}n(Al) = \frac{3}{2} \times 0.05 = 0.075 \;mol"

M(CuO) = 79.54 g/mol

m(CuO) "= 0.075 \times 79.54 = 5.965 \;g"

Answer: 5.96 g